# Toward Low-Altitude Airspace Management and UAV Operations: Requirements, Architecture and Enabling Technologies

### v1

@@
Guiyang Luo, Jinglin Li, Quan Yuan, Yijing Lin, and Ping Zhang are with the State Key Laboratory of Networking and Switching Technology, Beijing University of Posts and Telecommunications, Beijing 100876, China (e-mail: luoguiyang@bupt.edu.cn; jlli@bupt.edu.cn; yuanquan@bupt.edu.cn; yjlin@bupt.edu.cn; pzhang@bupt.edu.cn).Qixun Zhang, and Zhiyong Feng are with the Key Laboratory of Universal Wireless Communications, Ministry of Education, Beijing University of Posts and Telecommunications, Beijing 100876, China (e-mail: zhangqixun@bupt.edu.cn; fengzy@bupt.edu.cn).Hui Zhang is with the School of Computer and Information Tech- nology, Beijing Jiaotong University, Beijing 100876, China (e-mail: huizhang1@bjtu.edu.cn).Nan Cheng is with the Key State Laboratory of ISN and the School of Telecommunications Engineering, Xidian University, Xi’an 710071, China (e-mail: nancheng@xidian.edu.cn).

### v2

@@
Guiyang Luo, Jinglin Li, Quan Yuan, Yijing Lin, and Ping Zhang are with the State Key Laboratory of Networking and Switching Technology, Beijing University of Posts and Telecommunications, Beijing 100876, China (e-mail: luoguiyang@bupt.edu.cn; jlli@bupt.edu.cn; yuanquan@bupt.edu.cn; yjlin@bupt.edu.cn; pzhang@bupt.edu.cn) (corresponding author: Jinglin Li).Qixun Zhang, and Zhiyong Feng are with the Key Laboratory of Universal Wireless Communications, Ministry of Education, Beijing University of Posts and Telecommunications, Beijing 100876, China (e-mail: zhangqixun@bupt.edu.cn; fengzy@bupt.edu.cn).Hui Zhang is with the School of Computer and Information Tech- nology, Beijing Jiaotong University, Beijing 100876, China (e-mail: huizhang1@bjtu.edu.cn).Nan Cheng is with the Key State Laboratory of ISN and the School of Telecommunications Engineering, Xidian University, Xi’an 710071, China (e-mail: nancheng@xidian.edu.cn).